Solve the following differential equations changing the independent variables
X^2 d^2y/dx^2 - dy/dx - 4x^3 y= 8x^3 sinx^2 ,X>0.
1
Expert's answer
2020-03-20T17:51:41-0400

Solve


"x{d^2y \over dx^2}-{dy \over dx}-4x^3y=8x^3\sin(x^2),\ x>0"

Divide by "x"


"{d^2y \over dx^2}-{1 \over x}\cdot{dy \over dx}-4x^2y=8x^2\sin(x^2)"

Compare with


"y''+Py'+Qy=R"

We have


"P=-{1 \over x},Q=-4x^2,R=8x^2\sin(x^2)"

Chooze "z" such that "(dz\/dx)^2=4x^2" or "dz\/dx=2x" so that "z=x^2"


"{dy \over dx}={dy \over dz}\cdot{dz \over dx}=2x{dy \over dz}"

"{d^2y \over dx^2}=2{dy \over dz}+4x^2{d^2y \over dz^2}"

"2{dy \over dz}+4x^2{d^2y \over dz^2}-{1 \over x}(2x{dy \over dz})-4x^2y(z)=8x^2\sin(x^2)"

"{d^2y \over dz^2}-y(z)=2\sin z"

We will get


"(D_1^2-1)y=2\sin z"

"C.F.=c_1e^z+c_2e^{-z}"

"P.I.={1 \over D_1^2-1}(2\sin z)=2{1 \over -1^2-1}\cdot\sin z=-\sin z"

The required solution is


"y(z)=c_1e^z+c_2e^{-z}-\sin z"

"y=c_1e^{x^2}+c_2e^{-x^2}-\sin (x^2)"